# arXiv:2603.02009v1[math.AP] 02 Mar 2026

### v1

@@ 5.6. Step 5: The Final Bootstrap Argument
$X(T)\leq\underbrace{C_{S}E_{0}^{1/2}}_{\text{Initial data}}+\underbrace{\epsilon(N)}_{\text{High-freq. tail}}+\underbrace{C_{S}C_{a}N\sqrt{T}E_{0}^{1/2}}_{\text{Kelvin-Voigt penalty}}+\underbrace{C_{1}X(T)^{5}}_{\text{Nonlinearity}}.$ (32)
Here, the success of the argument hinges on a strict hierarchical choice of the parameters $N$ (the frequency threshold) and $T$ (the time horizon). We must proceed in the following precise order to avoid circular dependencies:
@@
    Fixing the Frequency $N$: The term $\epsilon(N)$ represents the energy of the high-frequency tail of the initial data. Since the initial data has finite energy, $\epsilon(N)\to 0$ as $N\to\infty$. We choose and fix a sufficiently large $N_{0}$ such that $\epsilon(N_{0})$ is arbitrarily small.
@@
    Fixing the Time $T$: With $N_{0}$ now fixed as a constant, the Kelvin-Voigt penalty term $C_{S}C_{a}N_{0}\sqrt{T}E_{0}^{1/2}$ solely depends on $T$. We can then choose the existence time $T>0$ sufficiently small so that this entire term is absorbed.
Once $N$ and $T$ are appropriately chosen, the linear terms and the small penalties can be consolidated into a single constant $K>0$. The inequality (32) then reduces to the standard polynomial form:
@@
To conclude that $X(T)$ is bounded, we invoke a standard bootstrap principle (or continuity argument). Consider the function $p(x)=x-C_{1}x^{5}$. For a sufficiently small constant $K$ (which can be ensured by shrinking the initial time interval), the equation $x-C_{1}x^{5}=K$ has two positive real roots, say $r_{1}<r_{2}$. This creates a "forbidden region" $(r_{1},r_{2})$ where the inequality $x-C_{1}x^{5}\leq K$ cannot be mathematically satisfied.
Since the mapping $t\mapsto X(t)=\|u\|_{L^{5}(0,t;L^{10}(\Omega))}$ is absolutely continuous in time, and it starts at $X(0)=0$ (which is strictly less than $r_{1}$), the value of $X(t)$ can never jump across the forbidden region $(r_{1},r_{2})$ to reach the larger values. Therefore, $X(T)$ must remain safely trapped in the bounded interval $0,r_{1}]$ for all times $t\in[0,T]$.
This establishes a uniform bound for the $L^{5}_{t}L^{10}_{x}$ norm that depends only on the initial data and the structural constants of the problem, but is completely independent of the truncation parameter $k$. This uniform bound is the ultimate key that allows us to pass to the limit as $k\to\infty$ and deduce the local existence of the strong solution.
@@ 5.8. Step 2: Uniform Higher Regularity Estimates via Strichartz
$\sup_{t\in[0,T]}\left(\|\nabla\partial_{t}u(t)\|_{2}^{2}+\|\Delta u(t)\|_{2}^{2}\right)\leq\mathcal{M}(E_{2}(0),E(0)),$ (47)
where $E_{2}(0)=\|\nabla u_{1}\|_{2}^{2}+\|\Delta u_{0}\|_{2}^{2}$, $E(0)$ is the standard first-order energy, and $\mathcal{M}$ is a constant strictly independent of any truncation parameter.
@@
Taking the $L^{2}(\Omega)$-inner product of the exact PDE for $u_{k}$ with the multiplier $\Delta\partial_{t}u_{k}$, we obtain the following identity:
@@
For the term involving the gradient of the damping coefficient, we employ the geometric assumption $|\nabla a|^{2}\leq ca(x)$. By writing $|\nabla a|=\frac{|\nabla a|}{\sqrt{a}}\sqrt{a}$ and applying Young’s inequality, we deduce:
@@
The first term on the right-hand side is absorbed by the strong dissipation on the left-hand side of (5.8), while the second term is bounded by $\frac{c}{2}E_{2}(t)$.
The most delicate part is the uniform estimate for the nonlinear term. Since $|f_{k}^{\prime}(s)|\leq 5|s|^{4}$ universally for all $k$, we can estimate the integral using the generalized Hölder inequality with exponents $(\frac{1}{3},\frac{1}{6},\frac{1}{2})$:
@@
Combining all the estimates back into (5.8), we arrive at the differential inequality:
@@
Notice that the integral inside the exponential is exactly the fourth power of the Strichartz norm $\|u_{k}\|_{L^{4}(0,t;L^{12}(\Omega))}^{4}$. As demonstrated in the subsequent sections, the Strichartz analysis guarantees that $\|u_{k}\|_{L^{4}L^{12}}\leq C_{S}$, where $C_{S}$ depends only on the initial energy $E(0)$ and is entirely independent of $k$.
@@
$\sup_{t\in[0,T]}\left(\|\nabla\partial_{t}u_{k}(t)\|_{2}^{2}+\|\Delta u_{k}(t)\|_{2}^{2}\right)\leq\mathcal{M}(E_{2}(0),E(0)),$ (55)
where $\mathcal{M}$ is a constant independent of $k$. This uniform bound is the cornerstone that allows us to extract a strongly convergent subsequence as $k\to\infty$, thereby proving the existence of strong solutions for the original critical quintic wave equation.
@@ 6.6. Removal of the truncation parameter $k$
Assuming failure of uniform bounds in $k$, we construct a renormalized sequence. Repeating the defect-measure argument shows again that $\mu\equiv 0$, producing the final contradiction. Indeed,
In previous works dealing with subcritical or Lipschitz non-linearities (e.g., [9]), the dependence on the truncation parameter $k$ is typically removed through a second contradiction argument, assuming $C_{k}\to\infty$ and applying the microlocal defect measure framework once again.
@@ 6.7. Conclusion
Appendix A Spectral Projectors and Bernstein Inequalities in Bounded Domains
In $\mathbb{R}^{n}$, the classical Littlewood-Paley decomposition relies heavily on the Fourier transform. For bounded domains $\Omega\subset\mathbb{R}^{3}$ with Dirichlet boundary conditions, the natural substitute is the spectral decomposition associated with the Laplace operator. In this appendix, we recall the definitions of the low and high-frequency spectral projectors and establish the corresponding Bernstein-type inequalities.
@@ A.1. Low-Frequency Projector and Bernstein’s Inequality
For a given frequency threshold $N>0$, we define the low-frequency projector $P_{\leq N}:L^{2}(\Omega)\to L^{2}(\Omega)$ by:
$P_{\leq N}u=\sum_{\sqrt{\lambda_{j}}\leq N}(u,\phi_{j})_{L^{2}}\phi_{j}.$ (71)
The operator $P_{\leq N}$ is an orthogonal projection in $L^{2}(\Omega)$, and hence it is self-adjoint. Moreover, it commutes with both $\Delta$ and $\partial_{t}$.
@@
$\|\nabla(P_{\leq N}u)\|_{L^{2}(\Omega)}\leq N\|u\|_{L^{2}(\Omega)}.$ (72)
@@
$\|\nabla(P_{\leq N}u)\|_{L^{2}(\Omega)}^{2}$ $=\sum_{\sqrt{\lambda_{j}}\leq N}\lambda_{j}|(u,\phi_{j})_{L^{2}}|^{2}$
$\leq\sum_{\sqrt{\lambda_{j}}\leq N}N^{2}|(u,\phi_{j})_{L^{2}}|^{2}$
$\leq N^{2}\sum_{j=1}^{\infty}|(u,\phi_{j})_{L^{2}}|^{2}=N^{2}\|u\|_{L^{2}(\Omega)}^{2}.$
@@
Remark A.2 (Orthogonality in the Energy Space and Regularity).
We implicitly assume that the boundary $\partial\Omega$ is sufficiently smooth (e.g., of class $C^{2}$) so that the domain of the Dirichlet Laplacian is fully characterized by $D(A)=H^{1}_{0}(\Omega)\cap H^{2}(\Omega)$. Furthermore, because the eigenfunctions $\phi_{j}$ are orthogonal with respect to the inner product in both $L^{2}(\Omega)$ and $H^{1}_{0}(\Omega)$, the projectors $P_{\leq N}$ and $P_{>N}$ are orthogonal projections in the energy space. In particular, for any $u\in H^{1}_{0}(\Omega)$, we have the exact exact decoupling of the gradient norm:
$\|\nabla u\|_{L^{2}(\Omega)}^{2}=\|\nabla(P_{\leq N}u)\|_{L^{2}(\Omega)}^{2}+\|\nabla(P_{>N}u)\|_{L^{2}(\Omega)}^{2}.$ (73)
This perfect splitting guarantees that separating the macroscopic wave $w_{N}$ from the microscopic fluctuation $v_{N}$ does not artificially inflate the total energy bounds during the a priori estimates.
@@ A.2. High-Frequency Projector and Poincaré-type Inequality
The high-frequency projector $P_{>N}:L^{2}(\Omega)\to L^{2}(\Omega)$ is defined naturally as the orthogonal complement:
$P_{>N}u=u-P_{\leq N}u=\sum_{\sqrt{\lambda_{j}}>N}(u,\phi_{j})_{L^{2}}\phi_{j}.$ (74)
@@
$\|P_{>N}u\|_{L^{2}(\Omega)}^{2}=\sum_{\sqrt{\lambda_{j}}>N}|(u,\phi_{j})_{L^{2}}|^{2}=\sum_{\sqrt{\lambda_{j}}>N}\frac{\lambda_{j}}{\lambda_{j}}|(u,\phi_{j})_{L^{2}}|^{2}.$
Since $\sqrt{\lambda_{j}}>N\implies\lambda_{j}>N^{2}$, we have $\frac{1}{\lambda_{j}}<\frac{1}{N^{2}}$ for all terms in the sum. Thus,
$\|P_{>N}u\|_{L^{2}(\Omega)}^{2}$ $\leq\frac{1}{N^{2}}\sum_{\sqrt{\lambda_{j}}>N}\lambda_{j}|(u,\phi_{j})_{L^{2}}|^{2}$

### v2

@@ 5.6. Step 5: The Final Bootstrap Argument
$X(T)\leq\underbrace{\|w_{N}^{lin}\|_{X_{T}}}_{\text{Low-freq linear}}+\underbrace{\epsilon(N)}_{\text{High-freq tail}}+\underbrace{C_{S}C_{a}N\sqrt{T}E_{0}^{1/2}}_{\text{Low-freq KV penalty}}+\underbrace{C_{a}\cdot T\cdot E_{0}^{1/2}}_{\text{High-freq KV commutator}}+\underbrace{C_{1}X(T)^{5}}_{\text{Nonlinearity}}.$ (46)
At first glance, closing a bootstrap argument with equation (46) might seem problematic for arbitrarily large initial data. The reader might rightfully wonder: how can we guarantee that the independent terms on the right-hand side are strictly small when the initial energy $E_{0}$ is large and the frequency cut-off $N$ increases the penalty?
This is where the underlying structure of our decomposition shines. The success of the argument hinges on a strict, non-circular hierarchical choice of the parameters $N$ (the frequency threshold) and $T$ (the time horizon). We must proceed in the following precise order:
@@
    Fixing the Frequency $N$ (Killing the Tail): The term $\epsilon(N)$ represents the Strichartz norm of the high-frequency linear evolution. Since the initial data has finite energy $E_{0}$, its spectral tail must vanish, meaning $\epsilon(N)\to 0$ as $N\to\infty$. We choose and strictly fix a sufficiently large $N_{0}$ such that $\epsilon(N_{0})$ is arbitrarily small.
@@
    Shrinking the Time $T$ (Taming the Regularized Data and the Penalties): With $N_{0}$ now permanently fixed, we analyze the remaining terms as $T\to 0$:
    • The low-frequency Kelvin-Voigt penalty $C_{S}C_{a}N_{0}\sqrt{T}E_{0}^{1/2}$ now solely depends on $\sqrt{T}$ and explicitly vanishes as $T\to 0$.
    • The high-frequency commutator source $C_{a}\cdot T\cdot E_{0}^{1/2}$ explicitly vanishes as $T\to 0$.
    • Most importantly, the low-frequency linear part $w_{N_{0}}^{lin}$ is now a fixed, highly regular function (since $N_{0}$ is fixed). By the absolute continuity of the Lebesgue integral, its norm over a shrinking time interval, $\|w_{N_{0}}^{lin}\|_{L^{5}(0,T;L^{10}(\Omega))}$, must strongly converge to zero as $T\to 0$, regardless of how large the initial energy $E_{0}$ originally was.
Once $N_{0}$ is fixed and $T$ is chosen sufficiently small, all the linear terms and penalties collectively collapse into a single, arbitrarily small constant $K>0$. The inequality (46) then gracefully reduces to the standard polynomial form:
@@
To conclude that $X(T)$ is bounded, we invoke a standard continuity argument (bootstrap principle). Consider the polynomial function $p(x)=x-C_{1}x^{5}$. For a sufficiently small constant $K$, the equation $x-C_{1}x^{5}=K$ has two positive real roots, say $r_{1}<r_{2}$. This creates a "forbidden region" $(r_{1},r_{2})$ where the algebraic inequality $x-C_{1}x^{5}\leq K$ cannot be mathematically satisfied.
Since the mapping $t\mapsto X(t)=\|u\|_{L^{5}(0,t;L^{10}(\Omega))}$ is continuous in time, and it originates at $X(0)=0$ (which is strictly less than the first root $r_{1}$), the value of $X(t)$ can never jump across the forbidden region to reach larger values. Therefore, $X(T)$ must remain safely trapped in the bounded interval $0,r_{1}]$ for all times $t\in[0,T]$.
This establishes a uniform bound for the critical $L^{5}_{t}L^{10}_{x}$ norm that depends only on the profile of the initial data and the structural constants of the problem, but is completely independent of the truncation parameter $k$. This uniform Strichartz bound is the fundamental key that allows us to safely pass to the limit as $k\to\infty$ in the truncated nonlinearity. Consequently, we rigorously deduce the local existence of a finite energy weak solution that additionally belongs to the critical Strichartz space. The higher regularity required to classify it as a strong solution will be extracted in the subsequent analysis.
@@ 5.8. Step 2: Uniform Higher Regularity Estimates via Strichartz
$\sup_{t\in[0,T]}\left(\|\nabla\partial_{t}u(t)\|_{2}^{2}+\|\Delta u(t)\|_{2}^{2}\right)\leq\mathcal{M}(E_{2}(0),E_{0}),$ (61)
where $E_{2}(0)=\|\nabla u_{1}\|_{2}^{2}+\|\Delta u_{0}\|_{2}^{2}$, $E_{0}$ is the standard first-order energy, and $\mathcal{M}$ is a constant strictly independent of any truncation parameter.
@@
Taking the $L^{2}(\Omega)$-inner product of the exact PDE for $u_{k}$ with the multiplier $\Delta\partial_{t}u_{k}$, and integrating by parts, we obtain the following identity:
@@
For the term involving the gradient of the damping coefficient, we employ the geometric assumption $|\nabla a|^{2}\leq ca(x)$. By writing $|\nabla a|=\frac{|\nabla a|}{\sqrt{a}}\sqrt{a}$ and applying the Cauchy-Schwarz and Young’s inequalities, we deduce:
@@
The first term on the right-hand side is gracefully absorbed by the strong dissipation on the left-hand side of (5.8), while the second term is bounded by $\frac{c}{2}E_{2}(t)$.
The most delicate part is the uniform estimate for the nonlinear term. Since $|f_{k}^{\prime}(s)|\leq 5|s|^{4}$ universally for all $k$, we can estimate the integral using the generalized Hölder inequality with the exact exponents $(\frac{1}{3},\frac{1}{6},\frac{1}{2})$:
@@
Combining all the estimates back into (5.8) and absorbing the fractional dissipation term, we arrive at the clean differential inequality:
@@
Notice that the integral inside the exponential is exactly the fourth power of the Strichartz norm $\|u_{k}\|_{L^{4}(0,t;L^{12}(\Omega))}^{4}$. As established in our prior Strichartz analysis, by bounding the total source term in $L^{1}(0,T;L^{2}(\Omega))$, the Strichartz estimates simultaneously control all critical pairs. In particular, we are guaranteed that $\|u_{k}\|_{L^{4}(0,T;L^{12}(\Omega))}\leq C_{S}$, where $C_{S}$ depends only on the initial energy $E_{0}$ and is strictly independent of $k$.
@@
$\sup_{t\in[0,T]}\left(\|\nabla\partial_{t}u_{k}(t)\|_{2}^{2}+\|\Delta u_{k}(t)\|_{2}^{2}\right)\leq\mathcal{M}(E_{2}(0),E_{0}),$ (69)
where $\mathcal{M}$ is a constant independent of $k$. This uniform bound is the rigorous cornerstone that allows us to extract a strongly convergent subsequence as $k\to\infty$, effectively passing to the limit and proving the existence of strong solutions for the original critical quintic wave equation. ∎
@@ 6.6. Removal of the truncation parameter $k$
Assuming failure of uniform bounds in $k$, we construct a renormalized sequence. Repeating the defect-measure argument shows again that $\mu\equiv 0$, producing the final contradiction. Indeed, in previous works dealing with subcritical or Lipschitz non-linearities (e.g., [9]), the dependence on the truncation parameter $k$ is typically removed through a second contradiction argument, assuming $C_{k}\to\infty$ and applying the microlocal defect measure framework once again.
@@ 6.7. Conclusion
Appendix A Smooth Spectral Projectors and Bernstein Inequalities
In $\mathbb{R}^{n}$, the classical Littlewood-Paley decomposition relies heavily on the Fourier transform. For bounded domains $\Omega\subset\mathbb{R}^{3}$ with Dirichlet boundary conditions, the natural substitute is the spectral decomposition associated with the Laplace operator. To ensure bounded commutators and rapid kernel decay, we must employ smooth spectral multipliers rather than sharp cutoffs. In this appendix, we define the smooth low and high-frequency spectral projectors via functional calculus and establish the corresponding Bernstein-type inequalities.
@@ A.1. Low-Frequency Projector and Bernstein’s Inequality
Let $\chi\in C_{c}^{\infty}(\mathbb{R})$ be an even, smooth bump function such that $0\leq\chi(s)\leq 1$ for all $s$, with $\chi(s)=1$ for $|s|\leq 1$ and $\chi(s)=0$ for $|s|\geq 2$. For a given frequency threshold $N>0$, we define the smooth low-frequency projector $P_{\leq N}:L^{2}(\Omega)\to L^{2}(\Omega)$ by:
$P_{\leq N}u=\sum_{j=1}^{\infty}\chi\left(\frac{\sqrt{\lambda_{j}}}{N}\right)(u,\phi_{j})_{L^{2}}\phi_{j}.$ (85)
Because $\chi$ is a bounded function, $P_{\leq N}$ is a bounded, self-adjoint operator in $L^{2}(\Omega)$. Moreover, it commutes with both $\Delta$ and $\partial_{t}$.
@@
$\|\nabla(P_{\leq N}u)\|_{L^{2}(\Omega)}\leq 2N\|u\|_{L^{2}(\Omega)}.$ (86)
@@
$\|\nabla(P_{\leq N}u)\|_{L^{2}(\Omega)}^{2}=\sum_{j=1}^{\infty}\lambda_{j}\left|\chi\left(\frac{\sqrt{\lambda_{j}}}{N}\right)\right|^{2}|(u,\phi_{j})_{L^{2}}|^{2}.$
Since the support of $\chi$ is contained in [2,2][-2,2], the terms in the sum are strictly zero whenever $\sqrt{\lambda_{j}}>2N$. For the non-zero terms, we have $\lambda_{j}\leq 4N^{2}$. Using the fact that $|\chi|\leq 1$, we bound the sum by:
$\|\nabla(P_{\leq N}u)\|_{L^{2}(\Omega)}^{2}$ $\leq\sum_{\sqrt{\lambda_{j}}\leq 2N}4N^{2}|(u,\phi_{j})_{L^{2}}|^{2}$
$\leq 4N^{2}\sum_{j=1}^{\infty}|(u,\phi_{j})_{L^{2}}|^{2}=4N^{2}\|u\|_{L^{2}(\Omega)}^{2}.$
@@
Remark A.2 (Energy Splitting).
Because we are using smooth multipliers instead of sharp cutoffs, $P_{\leq N}$ is not an idempotent projection (i.e., $P_{\leq N}^{2}\neq P_{\leq N}$). However, since $0\leq\chi(s)\leq 1$, the projectors $P_{\leq N}$ and $P_{>N}$ are uniformly bounded operators in the energy space. In particular, the algebraic identity $\chi(s)+(1-\chi(s))=1$ ensures $P_{\leq N}u+P_{>N}u=u$, guaranteeing that separating the macroscopic wave $w_{N}$ from the microscopic fluctuation $v_{N}$ does not artificially inflate the total bounds during the a priori estimates.
@@ A.2. High-Frequency Projector and Poincaré-type Inequality
The smooth high-frequency projector $P_{>N}:L^{2}(\Omega)\to L^{2}(\Omega)$ is defined as the complementary operator $P_{>N}=I-P_{\leq N}$, which translates to:
$P_{>N}u=\sum_{j=1}^{\infty}\left[1-\chi\left(\frac{\sqrt{\lambda_{j}}}{N}\right)\right](u,\phi_{j})_{L^{2}}\phi_{j}.$ (87)
@@
$\|P_{>N}u\|_{L^{2}(\Omega)}^{2}=\sum_{j=1}^{\infty}\left|1-\chi\left(\frac{\sqrt{\lambda_{j}}}{N}\right)\right|^{2}|(u,\phi_{j})_{L^{2}}|^{2}.$
By the definition of our cutoff function, $1-\chi(s)=0$ for all $|s|\leq 1$. Therefore, the terms in the sum are strictly zero for $\sqrt{\lambda_{j}}\leq N$. For the remaining terms where $\sqrt{\lambda_{j}}>N$, we have $\lambda_{j}>N^{2}$, which implies $1<\frac{\lambda_{j}}{N^{2}}$. Since $|1-\chi|\leq 1$, we can bound the sum by:
$\|P_{>N}u\|_{L^{2}(\Omega)}^{2}$ $\leq\sum_{\sqrt{\lambda_{j}}>N}1\cdot|(u,\phi_{j})_{L^{2}}|^{2}$
$\leq\sum_{\sqrt{\lambda_{j}}>N}\frac{\lambda_{j}}{N^{2}}|(u,\phi_{j})_{L^{2}}|^{2}$
@@
Appendix B Rigorous Commutator Estimates for the Damping Term
The objective of this appendix is to provide a rigorous proof of the commutator bound (37) utilized in the high-frequency analysis. Specifically, we must estimate the $L^{2}$-norm of the divergence of the commutator between the high-frequency projector $P_{>N}$ and the variable damping coefficient $a\in W^{1,\infty}(\Omega)$.
Let $f:=\nabla u_{t}\in L^{2}(\Omega)$. We aim to establish a uniform bound for:
$\left\|\operatorname{div}\Big([P_{>N},a(x)]f\Big)\right\|_{L^{2}(\Omega)}.$
Step 1: Reduction to Low Frequencies.
Handling the high-frequency projector directly is cumbersome. We exploit the identity $P_{>N}=I-P_{\leq N}$. Since the identity operator perfectly commutes with pointwise multiplication by $a(x)$, we have:
$P_{>N},a(x)]=[I-P_{\leq N},a(x)]=-[P_{\leq N},a(x)].$
Therefore, it suffices to estimate the divergence of the low-frequency commutator: $\operatorname{div}\big([P_{\leq N},a(x)]f\big)$.
Step 2: Integral Representation and Kernel Bounds.
By the functional calculus of the Dirichlet Laplacian, the smooth low-frequency projector $P_{\leq N}=\chi(\sqrt{-\Delta}/N)$ is an integral operator. Let $K_{N}(x,y)$ be its integral kernel. We write the action of the commutator $P_{\leq N},a]f$ explicitly as:
$\big([P_{\leq N},a]f\big)(x)$ $=P_{\leq N}(af)(x)-a(x)P_{\leq N}(f)(x)$
$=\int_{\Omega}K_{N}(x,y)a(y)f(y)\,dy-\int_{\Omega}K_{N}(x,y)a(x)f(y)\,dy$
$=\int_{\Omega}K_{N}(x,y)\Big(a(y)-a(x)\Big)f(y)\,dy.$
Because $\chi\in C_{c}^{\infty}(\mathbb{R})$ is an even, rapidly decaying function, the standard spectral cluster estimates for bounded domains (see, e.g., Sogge or Taylor) guarantee that the kernel $K_{N}(x,y)$ and its spatial derivatives satisfy the rapid decay bounds:
$|\nabla_{x}K_{N}(x,y)|\leq C_{M}\frac{N^{4}}{(1+N|x-y|)^{M}},$ (89)
for any integer $M>0$, where $C_{M}$ is a constant independent of $N$.
Step 3: Applying the Divergence Operator.
Next, we apply the spatial gradient $\nabla_{x}$ to the integral representation. By the product rule, the derivative acts on both the kernel $K_{N}$ and the coefficient $a(x)$:
$\nabla_{x}\big([P_{\leq N},a]f\big)(x)=\underbrace{\int_{\Omega}\nabla_{x}K_{N}(x,y)\big(a(y)-a(x)\big)f(y)\,dy}_{\mathcal{T}_{1}}-\underbrace{\int_{\Omega}K_{N}(x,y)\nabla a(x)f(y)\,dy}_{\mathcal{T}_{2}}.$
Step 4: Estimate for $\mathcal{T}_{2}$.
The second term can be rewritten simply as $\nabla a(x)\cdot(P_{\leq N}f)(x)$. Since $\nabla a\in L^{\infty}(\Omega)$ and the smooth spectral projections are uniformly bounded operators on $L^{2}(\Omega)$, we immediately obtain:
$\|\mathcal{T}_{2}\|_{L^{2}}\leq\|\nabla a\|_{L^{\infty}}\|P_{\leq N}f\|_{L^{2}}\leq C\|\nabla a\|_{L^{\infty}}\|f\|_{L^{2}}.$ (90)
Step 5: Estimate for $\mathcal{T}_{1}$ (The Cancellation Effect via Schur’s Test).
This term captures the core smoothing effect. Naively, the derivative $\nabla_{x}K_{N}$ costs a factor of $N^{4}$ (one $N$ more than the volume factor $N^{3}$). However, we control the difference $\big(a(y)-a(x)\big)$ using the Lipschitz continuity of $a\in W^{1,\infty}(\Omega)$ via the Mean Value Theorem:
$|a(y)-a(x)|\leq\|\nabla a\|_{L^{\infty}}|x-y|.$
Combining this with the kernel bound (89), we bound the modified kernel of $\mathcal{T}_{1}$, which we denote by $\tilde{K}_{N}(x,y)$, as follows:
$|\tilde{K}_{N}(x,y)|$ $=\Big|\nabla_{x}K_{N}(x,y)\Big|\cdot|a(y)-a(x)|$
$\leq C_{M}\|\nabla a\|_{L^{\infty}}\frac{N^{4}|x-y|}{(1+N|x-y|)^{M}}$
$=C_{M}\|\nabla a\|_{L^{\infty}}\cdot N^{3}\frac{N|x-y|}{(1+N|x-y|)^{M}}.$
Let us choose $M>4$. Because $\frac{s}{(1+s)^{M}}$ is bounded and integrable over $\mathbb{R}^{3}$, the integration of this modified kernel with respect to either $x$ or $y$ yields a uniform constant independent of $N$:
$\sup_{x\in\Omega}\int_{\Omega}|\tilde{K}_{N}(x,y)|\,dy\leq C\|\nabla a\|_{L^{\infty}},\quad\text{and}\quad\sup_{y\in\Omega}\int_{\Omega}|\tilde{K}_{N}(x,y)|\,dx\leq C\|\nabla a\|_{L^{\infty}}.$
By Schur’s Test for integral operators, this implies that the operator $\mathcal{T}_{1}$ maps $L^{2}(\Omega)$ to $L^{2}(\Omega)$ boundedly:
$\|\mathcal{T}_{1}\|_{L^{2}}\leq C\|\nabla a\|_{L^{\infty}}\|f\|_{L^{2}}.$ (91)
The diverging factor $N$ was perfectly cancelled by the spatial difference |xy||x-y|, absorbing the derivative without any asymptotic loss.
Conclusion.
Combining (90) and (91) via the triangle inequality, and recalling $f=\nabla u_{t}$, we conclude:
$\left\|\operatorname{div}\Big([P_{>N},a(x)]\nabla u_{t}\Big)\right\|_{L^{2}(\Omega)}\leq C\|\nabla a\|_{L^{\infty}}\|\nabla u_{t}\|_{L^{2}(\Omega)}\leq C_{a}E_{0}^{1/2}.$ (92)
This rigorously demonstrates that the divergence of the commutator is completely free of derivative loss, yielding a uniform bound independent of the frequency threshold $N$ in the bounded domain setting.
Remark B.1 (Pseudodifferential Calculus and Sogge’s Spectral Multipliers).
The rigorous kernel bound and Schur’s test detailed above confirm the heuristic expected from classical pseudodifferential calculus. In $\mathbb{R}^{3}$, the smooth projector $P_{\leq N}=\chi(\sqrt{-\Delta}/N)$ acts as a Fourier multiplier with symbol $\chi(|\xi|/N)$. The commutator $P_{\leq N},a(x)]$ generates a principal symbol given by the Poisson bracket:
$\sigma([P_{\leq N},a(x)])\sim\{\chi(|\xi|/N),a(x)\}=\nabla_{\xi}\left(\chi\left(\frac{|\xi|}{N}\right)\right)\cdot\nabla_{x}a(x)=\frac{1}{N}\nabla\chi\left(\frac{|\xi|}{N}\right)\cdot\nabla a(x).$
This explicitly reveals that the commutator is a regularizing operator of order 1-1. The crucial $1/N$ factor exactly cancels the spatial derivative coming from the divergence operator in the Kelvin-Voigt term, explaining the uniform $L^{2}$ bound independent of $N$.
Furthermore, because $\chi$ is a smooth multiplier, the celebrated spectral multiplier theorems for bounded domains by Sogge (see e.g., [28]) guarantee that the projectors $P_{\leq N}$ and $P_{>N}$ are uniformly bounded operators on $L^{p}(\Omega)$ for $1<p<\infty$. This deep fact ensures that separating the frequencies does not destroy the integrability required to close the critical Strichartz estimates in $L_{t}^{5}L_{x}^{10}(\Omega)$.